OpenStudy (anonymous):

Which of the following ordered pairs is a solution to the system of linear equations below? (Solutions: 4x + y = 19, 3x - y = 9) 1 and three sevenths, ten and three sevenths, (10,-21), (2,11),(4,3)

OpenStudy (amistre64):

when given options; just plug them in to see what works

OpenStudy (amistre64):

add the 2 equations together and you dont have to trial and error as much

OpenStudy (amistre64):

4x + y = 19 3x - y = 9 ----------- 7x + 0 = 28

OpenStudy (amistre64):

hmm, id go with the last option in this case :)